OpenStudy (anonymous):

Brain weight B as a function of body weight W in fish has been modeled by the power function B=.007W2/3, where B and W are measured in grams. A model for body weight as a function of body length L (measured in cm) is W=.12L2.53. If, over 10 million years, the average length of a certain species of fish evolved from 15cm to 20cm at a constant rate, how fast was the species' brain growing when the average length was 18cm? Round your answer to the nearest hundredth.

OpenStudy (kc_kennylau):

Just substitute L=18 into W=0.12*L^2.53, then into B=0.007*W^(2/3)

OpenStudy (anonymous):

just did and its not right :(

OpenStudy (anonymous):

don't I have to differentiate something? since this is a related rates problem :s